So 0<t<5 Using the Laplace transform, solve the initial value problem y + y = 3 t5 y(0) = 0. 9
Answer #1

Given y+y= ioltas { Glo)=0 3 ; t35 Note: Ylo) value not given we assume yo) =0 y ty luct) – ult-5)] + 3 u ult-5) y+y = o +By partial fraction decomposition 3 A. Bitc + 3(32+1) 1 12+1 For A > A= alain). 13=0 11 3 2 set 3 = 3 (0)²41 >> A =] 3 + mln3 2 د (41عدد و +212 + 2 + د3 د) 3 = (3+ B) Equating the Coefficients on both sides s from s2 - coefficient 3+B=0 = 2 = -3 froY(t) = 3244859} - 31]ses (3²+1 = 3u1-5) - 3 cos(4-5) Ultas) 11 3(1-сoslt-5)) ult-5) Answer is Y) = 3 [1-605(7-5)] ult-5) We (